# AraFinNews: Arabic Financial Summarisation with Domain-Adapted LLMs

### v1

@@ IV-A Model Selection and Comparative Setup
To benchmark Arabic financial summarisation effectively, we implemented several encoder–decoder transformer architectures within a unified text-to-text framework (Table IV). FinAraT5 was selected as the primary model due to its continued pretraining on large-scale Arabic financial documents, enabling it to capture domain-specific terminology, numerical reasoning, and stylistic conventions characteristic of financial reporting. For comparative evaluation, we included mT5 [21] as a multilingual baseline and AraT5 [7] as a general-purpose Arabic model. Together, these systems represent three distinct pretraining paradigms: multilingual generalisation, monolingual adaptation, and financial-domain specialisation.
@@
mT5-Base [21] Multilingual Encoder–Decoder (220M)
AraT5-Base [7] General Arabic Encoder–Decoder (220M)
FinAraT5-Base [37] Arabic Financial Encoder–Decoder (220M)
@@
All models adopt a text-to-text paradigm supporting multiple generation tasks through consistent input–output formatting. Each was fine-tuned on the AraFinNews corpus using identical hyperparameters to ensure comparability. Fine-tuning employed the Hugging Face Transformers library with the AdamW optimiser, a learning rate of $5\times 10^{-5}$ (with linear warm-up), an effective batch size of 32 achieved via gradient accumulation of four steps, and a maximum sequence length of 512 tokens. Evaluation focused on both linguistic quality and factual faithfulness to assess how domain adaptation influences summarisation performance.
@@ IV-B Arabic Financial Summarisation Pipeline
This study employs FinAraT5 [37], a domain-adapted generative language model fine-tuned primarily for Arabic financial headline generation and abstractive summarisation. Built upon the AraT5 architecture [7] and following the text-to-text paradigm established by T5 [20], FinAraT5 benefits from continued pretraining on large-scale Arabic financial corpora, including the AraFinNews dataset introduced in Section III. This adaptation enables the model to internalise the specialised terminology, discourse structures, and stylistic conventions characteristic of financial journalism.
The model is designed to generate concise, factually accurate, and stylistically coherent headlines that capture the numerical precision, entity references, and event-driven phrasing typical of Arabic financial reporting. Within our summarisation pipeline, FinAraT5 acts as the core component for abstractive headline generation, transforming full-length news articles into information-rich summaries aligned with the linguistic and stylistic expectations of Modern Standard Arabic financial communication.
@@
FinAraT5 was pretrained using the span-corruption denoising objective introduced in the original T5 framework [20], where contiguous spans of text are randomly masked and the model learns to reconstruct the missing content. This self-supervised approach enables large-scale learning without labelled data and effectively captures long-range contextual dependencies within text.
Following the methodology of AraT5 [7], FinAraT5 extends this objective through continued pretraining on raw Arabic financial documents, including large-scale news sources such as AraFinNews. This additional phase allows the model to internalise financial terminology, numerical expressions, and the discourse structures typical of professional economic reporting. As illustrated in Figure 4, the span-corruption objective provides an efficient mechanism for leveraging vast amounts of unannotated financial text while maintaining contextual richness and linguistic fidelity.
@@
Evaluation metrics ROUGE-1/2/L, METEOR, BERTScore, Faithfulness
@@ V-A Arabic Financial Headline Generation
This section presents experiments on the automatic generation of concise and informative Arabic financial headlines from full-length news articles in the AraFinNews dataset. The task represents a challenging form of abstractive summarisation requiring factual precision, linguistic coherence, and stylistic adherence to the conventions of Modern Standard Arabic (MSA) journalism. Unlike extractive summarisation, which selects sentences directly from the source, abstractive summarisation demands deep semantic interpretation and the generation of novel, well-formed text. In the financial domain, this challenge is heightened by the prevalence of numerical data, company names, and temporal expressions that must be faithfully reproduced.
We evaluate several pretrained transformer architectures—monolingual, multilingual, and domain-adapted—to assess how domain-specific pretraining affects Arabic financial summarisation. Our focus is on FinAraT5 [37], a Financial Arabic Large Language Model derived from AraT5 via continued pretraining on large-scale Arabic financial corpora, including data comparable in style and scope to AraFinNews. This configuration allows us to analyse whether specialised pretraining improves factual accuracy, semantic faithfulness, and stylistic precision.
@@
  • AraT5 (small, base) [7]: a monolingual Arabic text-to-text transformer trained on diverse general-domain corpora.
  • mT5 (small, base, large) [21]: a multilingual encoder–decoder model covering 101 languages, serving as a strong cross-lingual baseline.
@@
  • BERT2BERT [43]: a multilingual encoder–decoder model using mBERT for both encoding and decoding, included as a generic non-T5 generative baseline.
@@ V-B Evaluation
Model performance was evaluated using ROUGE [44], METEOR [45], and a suite of embedding-based metrics including BERTScore (BE), BARTScore (BA), FrugalScore [46], and BLEURT [47]. ROUGE and METEOR capture lexical and shallow semantic overlap, whereas BE, BA, and FrugalScore quantify embedding-based similarity, offering a complementary perspective on semantic faithfulness. BLEURT provides a learned measure of overall quality and coherence.
Results on the AraFinNews test split, presented in Table VI, exhibit trends consistent with those reported in the original FinAraT5 study [37]. FinAraT5 continues to outperform general-domain baselines, achieving gains of roughly +1.5–2 points across most evaluation metrics. These steady yet meaningful improvements suggest that financial-domain pretraining contributes to enhanced semantic accuracy and stylistic fluency, even when applied to a larger and more heterogeneous corpus such as AraFinNews.
Model BE BA Frugal-1 Frugal-2 BLEURT METEOR
@@ VI Conclusion
Experimental results showed that FinAraT5 consistently outperforms both monolingual and multilingual baselines across all automatic metrics, achieving improvements of up to +3.1 in ROUGE-L, +2.4 in METEOR, and +2.8% in FrugalScore over AraT5 and mT5. These gains highlight the effectiveness of domain-adaptive pretraining in producing summaries that are not only factually accurate and concise but also stylistically aligned with Arabic financial journalism. The analysis of attention distributions further demonstrated FinAraT5’s capacity to model fine-grained semantic and numerical relationships characteristic of financial language.
Overall, this work underscores the importance of domain adaptation for advancing Arabic NLP, particularly in specialised contexts where precision and factual grounding are essential. Future work will extend the AraFinNews framework to other Arabic business and policy domains, explore instruction-tuned and retrieval-augmented variants of FinAraT5, and investigate factual consistency and hallucination control through hybrid retrieval–generation architectures. Both the dataset and model will be released publicly to promote open, reproducible, and community-driven research in Arabic financial language processing.

### v2

@@ IV-A Model Selection and Comparative Setup
To benchmark Arabic financial summarisation effectively, we implemented several encoder–decoder transformer architectures within a unified text-to-text framework (Table IV). FinAraT5 was selected as the primary model due to its continued pretraining on large-scale Arabic financial documents, enabling it to capture domain-specific terminology, numerical reasoning, and stylistic conventions characteristic of financial reporting. For comparative evaluation, we included mT5 [xue-etal-2021-mt5] as a multilingual baseline and AraT5 [nagoudi_arat5_2022] as a general-purpose Arabic model. Together, these systems represent three distinct pretraining paradigms: multilingual generalisation, monolingual adaptation, and financial-domain specialisation.
@@
mT5-Base [xue-etal-2021-mt5] Multilingual Encoder–Decoder (220M)
AraT5-Base [nagoudi_arat5_2022] General Arabic Encoder–Decoder (220M)
FinAraT5-Base [zmandar2023finarat5] Arabic Financial Encoder–Decoder (220M)
@@
All models adopt a text-to-text paradigm supporting multiple generation tasks through consistent input–output formatting. Each was fine-tuned on the AraFinNews corpus using identical hyperparameters to ensure comparability. Fine-tuning employed the Hugging Face Transformers library with the AdamW optimiser, a learning rate of $5\times 10^{-5}$ (with linear warm-up), an effective batch size of 32 achieved via gradient accumulation of four steps, and a maximum sequence length of 512 tokens. Evaluation assesses how domain adaptation influences summarisation performance.
@@ IV-B Arabic Financial Summarisation Pipeline
This study employs FinAraT5 [zmandar2023finarat5], a domain-adapted generative language model fine-tuned primarily for Arabic financial headline generation and abstractive summarisation. Built upon the AraT5 architecture [nagoudi_arat5_2022] and following the text-to-text paradigm established by T5 [raffel_exploring_2020], FinAraT5 benefits from continued pretraining on large-scale Arabic financial corpora, including the AraFinNews dataset introduced in Section III. This adaptation enables the model to internalise the specialised terminology, discourse structures, and stylistic conventions characteristic of financial journalism.
Within our summarisation pipeline, FinAraT5 acts as the core component for abstractive headline generation, transforming full-length news articles into information-rich summaries aligned with the linguistic and stylistic expectations of Modern Standard Arabic financial communication.
@@
FinAraT5 was pretrained using the span-corruption denoising objective introduced in the original T5 framework [raffel_exploring_2020], where contiguous spans of text are randomly masked and the model learns to reconstruct the missing content. This self-supervised approach enables large-scale learning without labelled data and effectively captures long-range contextual dependencies within text.
Following the methodology of AraT5 [nagoudi_arat5_2022], FinAraT5 extends this objective through continued pretraining on raw Arabic financial documents, including large-scale news sources such as AraFinNews. This additional phase allows the model to internalise financial terminology, numerical expressions, and the discourse structures typical of professional economic reporting. As illustrated in Figure 4, the span-corruption objective provides an efficient mechanism for leveraging vast amounts of unannotated financial text while maintaining contextual richness and linguistic fidelity.
@@
Evaluation metrics ROUGE-1/2/L, METEOR, BARTScore, BERTScore, FrugalScore
@@ V-A Arabic Financial Headline Generation
This section presents experiments on the automatic generation of concise and informative Arabic financial headlines from full-length news articles in the AraFinNews dataset. The task represents a challenging form of abstractive summarisation requiring precision and coherence, and stylistic adherence to the conventions of Modern Standard Arabic (MSA) journalism. Unlike extractive summarisation, which selects sentences directly from the source, abstractive summarisation demands deep semantic interpretation and the generation of novel, well-formed text. In the financial domain, this challenge is heightened by the prevalence of numerical data, company names, and temporal expressions.
We evaluate several pretrained transformer architectures—monolingual, multilingual, and domain-adapted—to assess how domain-specific pretraining affects Arabic financial summarisation. Our focus is on FinAraT5 [zmandar2023finarat5], a Financial Arabic Large Language Model derived from AraT5 via continued pretraining on large-scale Arabic financial corpora, including data comparable in style and scope to AraFinNews.
@@
  • AraT5 (small, base) [nagoudi_arat5_2022]: a monolingual Arabic text-to-text transformer trained on diverse general-domain corpora.
  • mT5 (small, base, large) [xue-etal-2021-mt5]: a multilingual encoder–decoder model covering 101 languages, serving as a strong cross-lingual baseline.
@@
  • BERT2BERT [rothe-etal-2020-leveraging]: a multilingual encoder–decoder model using mBERT for both encoding and decoding, included as a generic non-T5 generative baseline.
@@ V-B Evaluation
Model performance was evaluated using ROUGE [lin_rouge_2004], METEOR [banerjee-lavie-2005-meteor], and a suite of embedding-based metrics including BERTScore (BertS), BARTScore (BartS), FrugalScore [kamal-eddine-etal-2022-frugalscore], and BLEURT [sellam2020bleurt]. ROUGE and METEOR capture lexical and shallow semantic overlap, whereas BERTScore, BARTScore, and FrugalScore quantify embedding-based similarity. BLEURT provides a learned measure of coherence.
Results on the AraFinNews test split, presented in Table VI, exhibit trends consistent with those reported in the original FinAraT5 study [zmandar2023finarat5]. FinAraT5 continues to outperform general-domain baselines, achieving gains of roughly +1.5–2 points across most evaluation metrics. These steady yet meaningful improvements suggest that financial-domain pretraining contributes to enhanced semantic accuracy and stylistic fluency, even when applied to a larger and more heterogeneous corpus such as AraFinNews.
Model BertS BartS Frugal-1 Frugal-2 BLEURT METEOR
@@ VI Conclusion
Experimental results showed that FinAraT5 consistently outperforms both monolingual and multilingual baselines across all automatic metrics, achieving improvements of up to +3.1 in ROUGE-L, +2.4 in METEOR, and +2.8% in FrugalScore over AraT5 and mT5. These gains highlight the effectiveness of domain-adaptive pretraining in producing summaries that are stylistically aligned with Arabic financial journalism. The analysis of attention distributions further demonstrated FinAraT5’s capacity to model fine-grained semantic and numerical relationships characteristic of financial language.